It has been known that a fixed dose of a certain drug results to an average increase of pulse rate by at least 12 beats per minute with a standard deviation of 5. A group of 20 patients given the same dose showed the following increases: 15, 12, 10, 8, 14, 15, 16, 11, 7, 13, 9, 10, 12, 11, 9, 10, 17, 14, 15, 7. Is there proof to show that this group has a lower average increase of pulse rate than the ones in general? Use 0.05 level of significance.


1
Expert's answer
2021-11-23T11:58:05-0500

"H_0:\mu=12" , group has not a lower average increase of pulse rate than the ones in general

"H_a:\mu<12" , group has a lower average increase of pulse rate than the ones in general


group mean:

"\overline{x}=\sum x_i/n=11.75"


group standard deviation:


"\sigma=\sqrt{\sum (x_i-\overline{x})^2/(n-1)}=3.02"


"t=\frac{\overline{x}-\mu}{\sigma/\sqrt n}=\frac{11.75-12}{5/\sqrt{20}}=-0.370"


"df=n-1=19"

from t-table for "\alpha=0.05" and lower one-sided test we get critical value:

"t_{crit}=1.729" for one-sided test


Since "|t|<t_{crit}" we accept the null hypothesis, that the group has not a lower average increase of pulse rate than the ones in general.